Project

General

Profile

log_template.txt

Stephen Kent, 09/06/2013 11:30 AM

 
1
*Observing Team*
2

    
3
_DES Observers:_ Rich Kron, Steve Kent, Ann Elliott, Niall MacCrann, Andres Plazas
4
_Telops:_ Hernan Tirado, Claudio Aguilera
5
_Remote Support:_
6

    
7
*Observing Plan*
8

    
9
# standard bias and dome flat sequence
10
# evening standard stars
11
# run ObsTac queue
12
# morning standard stars
13

    
14
*Accomplishments*
15

    
16
# Afternoon: zeros, dome flats (ugriz)
17
# Cloudy during evening twilight, so no standard-star fields.
18
# Waited for clearing and started at 22:40. Seeing much better tonight.  Variable cloudiness, but some intervals of good transparency. The following sequence is likely to have good seeing and reasonable transparency: 231046 - 231053. Also, the following sequence is likely to have reasonable seeing and good transparency:  231089 - 231096.
19
# Declared the following wide-field exposures to be "unaccepted by data management" in the mountaintop database:  229507 - 229596. 
20
# ObsTac attempted to re-observe SN-C3 but these exposures are bad (231116, 231117, 231118) due to thick clouds.  We marked these as "redo" and ObsTac re-loaded them onto the queue, but we had already stopped because of continuing clouds. 
21
# Again because of thick clouds, we also marked some wide-field exposures as "redo:"  231058, 231059, 231061, 231062, 231063, 231124.
22
# Overall: Including calibration exposures, this night extends from 230928 to 231134.  Science exposures extend from 231023 through 231124. No good SN exposures. No morning standard-star fields due to thick clouds after 02:40.
23

    
24
*Problems*
25

    
26
# See Problem Synopsis.
27

    
28
*Notes for Data Management*
29

    
30
# do not process 230987 - 230991 (20-second r flats)
31